# Visual-Inertial SLAM as Simple as A, B, VINS

### v2

@@ Abstract
We present AB-VINS, a different kind of visual-inertial SLAM system. Unlike most VINS systems which only use hand-crafted techniques, AB-VINS makes use of three different deep networks. Instead of estimating sparse feature positions, AB-VINS only estimates the scale and bias parameters (a and b) of monocular depth maps, as well as other terms to correct the depth using multi-view information which results in a compressed feature state. Despite being an optimization-based system, the main VIO thread of AB-VINS surpasses the efficiency of a state-of-the-art filter-based method while also providing dense depth. While state-of-the-art loop-closing SLAM systems have to relinearize a number of variables linear the number of keyframes, AB-VINS can perform loop closures while only affecting a constant number of variables. This is due to a novel data structure called the memory tree, in which the keyframe poses are defined relative to each other rather than all in one global frame, allowing for all but a few states to be fixed. AB-VINS is not as accurate as state-of-the-art VINS systems, but it is shown through careful experimentation to be more robust.
@@ I Introduction
Figure 1: A snapshot of the AB-VINS visualizer gives an overview of our system. On the bottom left is the current frame with feature tracks (red) and reprojected points (blue) overlayed. On the bottom center is the most recent keyframe depth map. On the bottom right is the code mask, which divides the images into different regions that are pushed and pulled to correct the depth according to multi-view information. In the main window at the top the memory tree can be seen, which is a novel data structure used to speed up pose graph optimization.
Visual-inertial simultaneous localization and mapping (SLAM) systems are crucial for augmented reality (AR) and virtual reality (VR). Leveraging the complimentary camera and inertial measurement unit (IMU) measurements, such systems output the high-rate metric, gravity-aligned poses (position and orientation) required for rendering, and bound the pose estimation error over time by remembering places they have been – performing loop closures to past states in order to prevent the user from drifting away in the virtual world. Besides the poses, dense depth is also useful for AR specifically. It is helpful for creating the illusion of occlusion with rendered objects, and it creates a solid surface for virtual characters to interact with.
Besides AR/VR, poses paired with dense depth are also needed for many robotic applications, such as obstacle avoidance and path planning. Despite the need for dense depth, very few VINS systems actually provide it. Additionally, despite recent advances in deep learning for a multitude of related applications, very few VINS systems have actually utilized deep networks. The proposed AB-VINS on the other hand provides high-rate gravity-aligned poses along with dense depth, and uses three different deep networks to its advantage. Visual-inertial SLAM needs to be as efficient as possible, being able to quickly close loops while running on some of the smallest devices – even as small as a bee robot in future applications. AB-VINS contains, to the best of our knowledge, the most efficient pose graph SLAM solution ever proposed thanks to the novel memory tree, which brings us one step closer to achieving such goals. An overview of AB-VINS can be seen in Fig. 1. Our main contributions are:
@@
  • A compact feature representation called AB features is presented, where an arbitrary number of feature positions are parameterized by $a$ and $b$, the scale and bias of a monocular depth map, as well as other terms to correct the depth with multi-view information.
@@
The rest of the paper is organized as follows: After the complete review of related work in Sec. II, we present in detail the proposed AB-VINS in Sec. III. In Sec. IV AB-VINS is thoroughly evaluated for accuracy, efficiency, and robustness. Finally, in Sec. V, closing remarks are made.
Figure 2: The proposed AB-VINS system.
@@ II Related Work
The proposed AB-VINS is related to both visual-inertial SLAM and dense SLAM systems. Because of the introduction of the memory tree, AB-VINS is also related to efficient pose graph optimization methods.
II-A Visual-Inertial SLAM Systems
@@
A multitude of visual-inertial SLAM systems have been presented in recent years. The most similar one in nature to AB-VINS is VI-ORB-SLAM [1], which is incorporated into ORB-SLAM3 [2]. Similar to AB-VINS, VI-ORB-SLAM has three modules: tracking, local mapping, and mapping. Just like AB-VINS, it is an optimization-based system, but despite using one more CPU thread than AB-VINS, it does not provide dense depth. In fact, none of the systems mentioned in this section provide dense depth. VINS-Mono [3], later called VINS-Fusion, similarly to AB-VINS has two threads – one for visual-inertial odometry (VIO) and one for mapping. Another visual-inertial SLAM system, ROVIOLI [4], similarly has a VIO front-end and pose graph back-end, and supports multi-session mapping and localization. BASALT [5] attempts to solve the problem of degraded IMU information for keyframes, and solves loop closures in a global bundle adjustment (BA), which may be computationally expensive. OKVIS2 [6] is a recent visual-inertial SLAM system which utilizes a semantic segmentation network to ignore features that should not be tracked (such as on clouds). The main drawback of all of these methods is that the number of variables affected by the global optimization grows linearly with the number of keyframes, while AB-VINS can solve loop closures while only relinearizing a constant number of variables.
@@
Unlike visual-inertial SLAM systems, pure visual-inertial odometry (VIO) methods do not have a memory about places, and thus the system steadily drifts over time – but is highly accurate locally. Many VIO systems have been proposed. The multi-state constraint Kalman filter (MSCKF) [7] is one of the first examples of a pure VIO system. MSCKF avoids estimating the structure in the extended Kalman filter (EKF) framework, which can be expensive with more than just a few features, by applying a nullspace projection trick to eliminate the dependency on the feature states. MSCKF has been adapted and improved many times by adding small amounts of structure to the state [8], adding online calibration and improving consistency [9], and utilizing the efficient sparse square root inverse covariance EKF form [10]. Robocentric VIO (R-VIO) [11], which is also based on the EKF, showed that estimating the poses in a local robocentric frame (the oldest frame in the sliding window) rather than in the global as usual has better observability properties. R-VIO partially inspired the anchored local mapping window of AB-VINS (discussed in Sec. III-C3), however instead of defining the anchored frame in the last frame of the sliding window, we define it in the gravity-aligned frame centered at the oldest frame in the window so that the local gravity direction does not have to be estimated. Another popular type of VIO system is based on nonlinear optimization. A popular method is to marginalize states that are removed from the optimization in order to maintain a large covariance matrix for additional prior cost terms [12, 6, 3, 5]. However, the computational cost of marginalization is high in optimization-based systems compared to EKF-based systems due to the cost of calculating the marginal covariance, which is the main reason we choose to ommit marginalization in AB-VINS and instead opt to simply fix the oldest pose in the sliding window similarly to ORB-SLAM.
II-B Dense SLAM Systems
@@
Dense VINS systems provide both high-rate gravity-aligned poses and dense depth. Kimera [13] provides dense depth via Delaunay triangulation on the sparse VIO points, however this is an oversimplification of the scene and can not capture small details of the structure. CodeVIO [14] is a dense VIO system which, similarly to AB-VINS, estimates correction terms for the depth within the state, but the correction terms (codes) are learned from an autoencoder network. The codes in AB-VINS are more hand-crafted, but can more easily push or pull entire objects with only a few feature measurements. DiT-SLAM [15] showed that the code dimension can be as low as 8. In AB-VINS on the other hand, the code dimension is 4. SimpleMapping [16] utilizes a deep multi-view stereo (MVS) network to provide dense depth, but does not utilize monocular priors to help the system when parallax is not available.
@@
Purely visual dense SLAM systems have also been proposed. The scale can not be estimated using only classical methods for monocular visual SLAM. DTAM [17] is the first example of such a system, which estimates the pose via dense image alignment, and creates dense depth maps with classical MVS. TANDEM [18], which pairs an MVS network with monocular visual odometry, and receives localization measurements from the TSDF map it maintains from fusing the dense depth. Some visual SLAM systems, like CNN-SLAM [19], BA-Net [20], CodeSLAM [21], and Deep Factors [22] can estimate metric scale by training a monocular depth network with metric depth supervision. However, training a network in such a way does not benefit from the robustness that eliminating the scale and bias from monocular depth as in MiDaS [23] provides. AB-VINS can estimate metric scale because of the IMU, and still benefits from the robustness of MiDaS.
II-C Efficient Pose Graph Optimization Methods
@@
One of the first efficient solutions to pose graph optimization is sparse pose adjustment (SPA) [24], which utilizes the sparse information matrix structure in a Levenberg–Marquardt optimization to its advantage. Another option is to utilize the square-root information form as in iSAM [25] and iSAM2 [26]. iSAM2 introduced a data structure called the Bayes tree, which is used to reduce the number of times variables need to be relinearized by using the tree to identify which parts of the graph are affected by the measurements – making optimization highly-efficient. The proposed memory tree is not related to the Bayes tree since the keyframe poses are defined in a relative frame rather than the global. SPA, iSAM, and iSAM2, as well as most other factor graph optimizations ever proposed, define all of the poses in the global frame, which requires at least adjusting all of the poses in-between two loop keyframes. Most similar to the proposed memory tree data structure is the relative method [27], where a keyframe’s pose is defined relative to the previous one’s frame of reference. While the relative method [27] was only applied to an incremental BA and not pose graphs, it could theoretically be used for pose graphs as well – just like the memory tree could be applied to an incremental BA. Just like the memory tree, the relative method allows for the majority of the poses to be fixed, however the proposed memory tree has better algorithmic properties such as theoretically smaller maximum state size and time to compute the global pose as well as optimization with better worst-case complexity.
III AB-VINS
In this section, the proposed AB-VINS system is detailed in full. A diagram of AB-VINS can be seen in Fig. 2.
III-A Initialization
@@
While some recent methods have utilized monocular depth to improve visual-inertial initialization [28, 29], those initialization procedures still wait some time to initialize, and require at least a small amount of motion. In AB-VINS, we adopt a much simpler initializer. We assume that the IMU biases are known well enough to integrate, which can easily be obtained offline by letting the device sit still. We first estimate the gravity direction by running Gram–Schmidt on the first few accelerometer measurements. While this does make a static assumption, we found that it works well except in the most dynamic scenarios, which may not be typical for AR/VR headsets or many robotics applications (e.g., wheeled robots or MAV inspection) anyway. Then, we initialize the gravity-aligned map with the first frame’s monocular depth using an arbitrary scale $a_{0}$, $b_{0}$. Finally, the velocity is set to zero, and tracking to the static map commences. Thus, we can typically initialize using just the first frame. No motion is required to initialize, but some excitation is needed to estimate the scale. However, we argue that with very little motion (sitting mostly or completely still) scale is not as important as it is with large motions.
@@ III-B Tracking to Static Map
For an incoming frame at time $t_{k}$, feature tracks are first obtained using the efficient Kenade-Lucas-Tomasi (KLT) tracking algorithm [30] on FAST corner points [31]. Then, the current pose is predicted from preintegrated measurements between the last frame and the current. After that, tracking to the static local map occurs. We track 123 feature points by default.
A visualization of tracking to a static map can be seen in Fig. 3. First the 6-DoF pose, represented as a JPL quaternion [32] ${}_{A}^{I_{k}}\bar{q}$ and 3D position vector ${}^{A}\mathbf{p}_{I_{k}}$, are estimated with only visual measurements – reprojecting points from the static local map  [^1] Note that since the local map is scaled and gravity-aligned, the pose output of the vision-only optimization is scaled and gravity-aligned since it is equivalent to the perspective $n$ point (PnP) problem with a good initial guess from the IMU.. $\{A\}$ is the gravity-aligned anchor frame of the local mapping window (discussed in the next section), and $\{I_{k}\}$ is the IMU frame at time $t_{k}$. Formally, the vision-only optimization attempts to minimize a cost function $\mathcal{C}_{vo}$ defined as
where $\mathcal{V}_{k}$ is the index set of all 3D points ${}^{A}\mathbf{p}_{f_{\ell}}$ visible in the current frame, $\lambda_{\ell}$ is the edge weight proposed by [29], which down-weights the cost for residuals corresponding to depth points that are near the image edge or have a high depth edge value. The edge images correspond to the depth map that the 3D point ${}^{A}\mathbf{p}_{f_{\ell}}$ originates from. The function $\rho_{c}$ is the robust Cauchy loss function, $\bm{\Sigma}_{v}$ is the diagonal noise covariance matrix for the residual (set to identity in our implementation), and the residual $\mathbf{r}^{rs}_{\ell k}$ is defined as
where $h_{rs}$ projects the point ${}^{A}\mathbf{p}_{f_{\ell}}$ into the current frame using the current pose estimate, $\mathbf{u}_{\ell k}$ is the raw pixel coordinate of the current feature track location for the $\ell$th 3D point, and ${}^{C}_{I}\bar{q},{}^{C}\mathbf{p}_{I},\bm{\zeta}$ are the camera extrinsic and intrinsic calibration parameters from the local mapper. Note that the 3D point and calibration are fixed for this optimization. The optimization is run up to four times, and after each step each residual is run through a $\chi^{2}$ check. If the residual fails the $\chi^{2}$ check, it is removed from the next round of optimization. At most 50% of residuals can be removed this way. This process robustifies against bad feature tracks or noisy 3D points.
After the vision-only optimization, the velocity ${}^{A}\mathbf{v}_{I_{k}}$ and IMU biases $\mathbf{b}_{g,k}$, $\mathbf{b}_{a,k}$ are estimated in an inertial-only optimization (only preintegration measurements) with poses of the previous keyframe and current frame fixed. For the preintegration, the angular velocity is integrated using the quaternion integrator [32] and the acceleration is double integrated using fourth order Runge-Kutta. In the case where the IMU measurements (angular velocity or linear acceleration) exceed a saturation threshold, the corresponding measurement noise is inflated to robustify the system against IMU saturation. In particular, the inertial-only optimization tries to minimize the cost function $\mathcal{C}_{io}$ where
@@
is the preintegration [33, 34, 35] cost function between the previous keyframe at time $t_{i}$ and the current frame at time $t_{k}$. We have the state $\hat{\bar{\mathbf{x}}}_{k}=\begin{bmatrix}{}_{A}^{I_{k}}{\bar{q}}^{\top}&{}^{A}{\mathbf{p}}_{I_{k}}^{\top}&{}^{A}\hat{\mathbf{v}}_{I_{k}}^{\top}&\hat{\mathbf{b}}_{g,k}^{\top}&\hat{\mathbf{b}}_{a,k}^{\top}\end{bmatrix}^{\top}$ which is an inertial state (note that the pose is fixed), and similarly for the completely fixed keyframe state $\mathbf{x}_{i}$. The terms ${}^{i}\bm{\alpha}_{k}$ and ${}^{i}\bm{\beta}_{k}$ are the preintegrated measurements between $t_{i}$ and $t_{k}$ defined as
@@
$\mathbf{Q}_{ik}$ is the covariance matrix for the preintegration residual, which can be inflated as discussed if there is IMU saturation detected.
In AB-VINS, we choose a tracking optimization that decouples the visual and inertial measurements. This is because we found that tightly coupling the visual and inertial measurements in the tracking optimization (e.g., as in [1]) can sometimes pull the solution away from one that satisfies the reprojection error, as seen in Fig. 4, which creates wobblier pose output than desired for many applications. The problem is exacerbated when tracking a small number of features as in our system – since the vision terms can not outweigh the inertial – and the noisiness of the monocular depth does not help either. We emphasize that our decoupled approach is only possible with nonlinear optimization and not with a filter, as filters naturally tightly couple the measurements by using inertial measurements for propagation and visual ones for update – which adds priors from both measurements to the pose states.

### v3

@@ Abstract
We present AB-VINS, a different kind of visual-inertial SLAM system. Unlike most popular VINS methods which only use hand-crafted techniques, AB-VINS makes use of three different deep neural networks. Instead of estimating sparse feature positions, AB-VINS only estimates the scale and bias parameters ($\mathbf{a}$ and $\mathbf{b}$) of monocular depth maps, as well as other terms to correct the depth using multi-view information, which results in a compressed feature state. Despite being an optimization-based system, the front-end motion tracking thread of AB-VINS surpasses the efficiency of a state-of-the-art filtering-based method while also providing dense depth. When performing loop closures, standard keyframe-based SLAM systems need to relinearize a number of variables which is linear with respect to the number of keyframes. In contrast, the proposed AB-VINS can incorporate loop closures while only affecting a constant number of variables. This is thanks to a novel data structure called the memory tree, where keyframe poses are defined relative to each other rather than all in one global frame, allowing for all but a few states to be fixed. While AB-VINS might not be as accurate as state-of-the-art VINS algorithms, it is shown to be more robust.
@@ I Introduction
Visual-inertial SLAM (or VINS) is essential for autonomous robots, augmented reality (AR) and virtual reality (VR) or extended reality (XR) [1]. Leveraging the complimentary camera and IMU measurements, such systems output high-rate, metric, gravity-aligned poses (orientations and positions) required for XR rendering, while bounding the pose estimation error over time by performing loop closures to past states in order to prevent the user from drifting away in the virtual world. Besides the poses, dense depth is useful, which, for instance, helps to create the illusion of occlusion with rendered object and solid surfaces for virtual characters to interact with. In robotic applications, poses paired with dense depth are also needed, for example, for obstacle avoidance and path planning.
Despite the need for dense depth, few VINS solutions are able to provide it. Additionally, despite recent advances in deep learning for a multitude of related applications [2, 3, 4], very few VINS methods have utilized it. AB-VINS on the other hand provides high-rate gravity-aligned poses along with dense depth, and uses three different deep networks to its advantage. Visual-inertial SLAM often needs to be as efficient as possible, being able to quickly close loops while running on some of the smallest devices – even as small as a bee robot in future applications. The proposed AB-VINS contains, to the best of our knowledge, the most efficient pose graph SLAM solution thanks to the novel memory tree data structure, which brings us one step closer to achieving such goals.
In particular, Fig. 1 depicts the overall system of the proposed AB-VINS which has the following main contributions:
@@
  • A compact feature representation called AB features is presented, where an arbitrary number of feature positions are parameterized by $a$ and $b$, the scale and bias of a monocular depth map, as well as other optional terms to correct the depth with multi-view information.
@@
Figure 1: Visualizing the proposed AB-VINS: On the bottom left is the current frame with feature tracks (red) and reprojected points (blue) overlayed. On the bottom center is the most recent keyframe depth map. On the bottom right is the code mask, which divides the images into different regions that are pushed and pulled to correct the depth according to multi-view information. In the main window at the top the memory tree can be seen, which is a novel data structure used to speed up pose graph optimization.
The rest of the paper is organized as follows: After reviewing the related work in Sec. II, we present in detail the proposed AB-VINS in Sec. III and IV. In Sec. V AB-VINS is thoroughly evaluated for accuracy, efficiency, and robustness. Finally, in Sec. VI, closing remarks are made.
@@ II Related Work
In this section, we provide a complete review of the literature that is closely related to the proposed AB-VINS, including visual-inertial SLAM, dense SLAM and efficient pose graph optimization methods.
II-A Visual-Inertial SLAM
@@
A multitude of visual-inertial SLAM systems have been developed in recent years [1]. The most similar one in nature to our AB-VINS is VI-ORB-SLAM [5], which is incorporated into ORB-SLAM3 [6]. Similar to AB-VINS, VI-ORB-SLAM is an optimization-based system and has three modules: tracking, local mapping, and mapping. Despite using one more CPU thread than AB-VINS, it does not provide dense depth (and actually none of the systems mentioned below do so). VINS-Mono [7], and later VINS-Fusion [^1] Available at: https://github.com/HKUST-Aerial-Robotics/VINS-Fusion, similarly to AB-VINS have two threads, one for visual-inertial odometry (VIO) and one for mapping, although VIO can be performed on two threads if desired. Another visual-inertial SLAM system, ROVIOLI [8], also has a VIO front-end and pose graph back-end, and supports multi-session mapping and localization. The recent work [9] similarly has a VIO front-end and optimization-based back-end, and aims at producing state estimates with covariances that are consistent with the true estimation error. BASALT [10] attempts to solve the problem of degraded IMU information for keyframes, and solves loop closures in a global bundle adjustment (BA), which may be computationally expensive. OKVIS2 [11] utilizes a semantic segmentation network to ignore features that should not be tracked (such as on clouds). The main drawback of all of these methods is that the number of variables affected by the global optimization grows linearly with the number of keyframes, while AB-VINS can solve loop closures while only relinearizing a constant number of variables.
@@
Unlike visual-inertial SLAM, pure VIO methods do not have a memory about places, and thus the system steadily drifts over time – but is highly accurate locally. Many VIO systems have been proposed [1]. The multi-state constraint Kalman filter (MSCKF) [12] is one of the first such examples. MSCKF avoids estimating the feature positions in the EKF framework, which can be expensive with more than just a few features, by applying the nullspace projection to eliminate the dependency on the feature states. Over the years, MSCKF has been adapted and improved, for example, by adding small amounts of structure to the state [13], adding online calibration and improving consistency [14], and utilizing the efficient sparse square root inverse covariance filter [15]. Robocentric VIO (R-VIO) [16], which is also based on the EKF, showed that estimating the poses in a local robocentric frame (the oldest frame in the sliding window) rather than in the global as usual has better observability properties. R-VIO partially inspired the anchored local mapping window of AB-VINS (discussed in Sec. III-C3), however instead of defining the anchored frame in the last frame of the sliding window, we define it in the gravity-aligned frame centered at the oldest frame in the window so that the local gravity direction does not have to be estimated. Another popular type of VIO system is based on nonlinear optimization, which often marginalizes states that are removed from the optimization in order to maintain a large covariance matrix for additional prior cost terms [17, 11, 7, 10]. However, the computational cost of marginalization is high in optimization-based systems compared to EKF if calculating the marginal covariance [18], motivating us to choose to ommit marginalization in AB-VINS and instead opt to simply fix the oldest pose in the sliding window similarly to ORB-SLAM.
II-B Dense SLAM
@@
Dense VINS can provide both high-rate gravity-aligned poses and dense depth. Kimera [19] provides dense depth via Delaunay triangulation on the sparse VIO points, however this is an oversimplification of the scene and can not capture small details of the structure. CodeVIO [20] is a dense VIO system which, similarly to AB-VINS, estimates correction terms for the depth within the state, but the correction terms (codes) are learned from an autoencoder network. The codes in AB-VINS are more hand-crafted, but can more easily push or pull entire objects with only a few feature measurements. DiT-SLAM [21] showed that the code dimension can be as low as 8. In AB-VINS, on the other hand, the code dimension is 4. SimpleMapping [22] utilizes a deep multi-view stereo (MVS) network to provide dense depth, but does not utilize monocular priors to help the system when parallax is not available.
@@
Pure visual dense SLAM systems have also been proposed. The scale can not be estimated using only classical methods for monocular visual SLAM. DTAM [23] is the first example of such a system, which estimates the pose via dense image alignment, and creates dense depth maps with classical MVS. TANDEM [24], which pairs an MVS network with monocular visual odometry, and receives localization measurements from the TSDF map it maintains from fusing the dense depth. Some visual SLAM systems, like CNN-SLAM [25], BA-Net [26], CodeSLAM [27], and Deep Factors [28] can estimate metric scale by training a monocular depth network with metric depth supervision. However, training a network in such a way does not benefit from the robustness that eliminating the scale and bias from monocular depth as in MiDaS [2] provides. AB-VINS can estimate metric scale because of the IMU, and still benefits from the robustness of MiDaS.
Figure 2: System diagram of the proposed AB-VINS.
II-C Efficient Pose Graph Optimization
@@
One of the first efficient solutions to pose graph optimization is the sparse pose adjustment (SPA) [29], which utilizes the sparse information matrix structure in a Levenberg–Marquardt (LM) optimization to its advantage. Another option is to utilize the square-root information form as in iSAM [30] and iSAM2 [31]. iSAM2 introduced a data structure called Bayes tree, which is used to reduce the number of times variables need to be relinearized by using the tree to identify which parts of the graph are affected by the measurements – thus making optimization highly-efficient. The proposed memory tree is not related to the Bayes tree since the keyframe poses are defined in a relative frame rather than the global. SPA, iSAM, and iSAM2, as well as most other factor graph optimization methods ever proposed, define all of the poses in the global frame, which requires at least adjusting all of the poses in-between two loop keyframes. Most similar to the proposed memory tree data structure is the relative method [32], where a keyframe’s pose is defined relative to the previous one’s frame of reference. While the relative method [32] was only applied to an incremental BA and not pose graphs, it could theoretically be used for pose graphs as well – just like the memory tree could be applied to an incremental BA. The relative method allows for the majority of the poses to be fixed, but the proposed memory tree has better algorithmic properties such as theoretically smaller maximum state size and time to compute the global pose as well as optimization with better worst-case complexity.
III Tracking and Local Mapping
We now explain the proposed AB-VINS in detail. Overall, as shown in Fig. 2, the system is composed of two threads: VIO and mapping. For every frame, the main VIO thread tracks the pose to the current static local map. When a frame is selected to be a keyframe, a depth map and code mask are generated and then sent to the anchored local mapping module, which optimizes the local structure and pose with both visual and inertial measurements in a sliding window fashion along with online calibration. Once a keyframe is popped off of the local mapping window, it is combined with the outputs of two more neural networks (i.e., one for global descriptors and the other for local keypoint descriptors) before being sent to the mapper, which communicates with the proposed memory tree and is responsible for finding loop closures. In the following, we will first focus on the main VIO thread, while the mapper thread will be discussed in the next section.
III-A Fast Initialization
@@
While some recent methods have utilized learned monocular depth to improve visual-inertial initialization [33, 34, 35], they still wait some time to initialize and require at least a small amount of motion. In AB-VINS, we adopt a much simpler initializer by assuming the IMU biases are known well enough to integrate (which can be obtained offline by letting the device sit still). We first estimate the gravity direction by running Gram–Schmidt on the first few accelerometer measurements. While this does make a static assumption, we found that it works well except in the most dynamic scenarios, which may not be typical for AR/VR headsets or mobile robotic applications (e.g., indoor wheeled robots). Then, we initialize the gravity-aligned map with the first frame’s monocular depth using an arbitrary scale and bias (i.e., $a_{0}$ and $b_{0}$). Finally, the velocity is set to zero, and tracking to the static map commences. Thus, we can typically initialize using just the first frame. No motion is required to initialize, but some excitation is needed to estimate the metric scale. However, we argue that with very little motion (sitting mostly or completely still) scale is not as important as it is with large motions.
@@ III-B Tracking to Static Map
For an incoming frame at time $t_{k}$, feature tracks are first obtained using the efficient Kenade-Lucas-Tomasi (KLT) tracking algorithm [36] on FAST corner points [37]. Then, the current pose is predicted from preintegrated measurements between the last frame and the current. After that, tracking to the static local map occurs. A visualization of this process can be seen in Fig. 3.
First, the 6-DoF pose, represented by the JPL quaternion [38] ${}_{A}^{I_{k}}\bar{q}$ and 3D position vector ${}^{A}\mathbf{p}_{I_{k}}$, are estimated with only visual measurements by reprojecting points from the static local map. [^2] Note that since the local map is scaled and gravity-aligned, the pose output of the vision-only optimization is scaled and gravity-aligned since it is equivalent to the PnP problem with a good initial guess from the IMU. Here, $\{A\}$ is the gravity-aligned anchor frame of the local mapping window and $\{I_{k}\}$ is the IMU frame at time $t_{k}$. Formally, the vision-only optimization attempts to minimize a cost function $\mathcal{C}_{vo}$ given by:
where $\mathcal{V}_{k}$ is the index set of all 3D points ${}^{A}\mathbf{p}_{f_{\ell}}$ tracked in the current frame, $\lambda_{\ell}$ is the edge weight proposed by [35], which down-weights the cost for residuals corresponding to depth points that are near the image edge or have a high depth edge value. The edge images correspond to the depth map that the 3D point ${}^{A}\mathbf{p}_{f_{\ell}}$ originates from. The function $\rho_{c}$ is the robust Cauchy loss function, and $\bm{\Sigma}_{v}$ is the diagonal noise covariance matrix for the residual $\mathbf{r}^{rs}_{\ell k}$ which is defined as:
where $h_{rs}$ projects the point ${}^{A}\mathbf{p}_{f_{\ell}}$ into the current frame using the current pose estimate, $\mathbf{u}_{\ell k}$ is the raw pixel coordinate of the current feature track location for the $\ell$-th 3D point, and ${}^{C}_{I}\bar{q},{}^{C}\mathbf{p}_{I},\bm{\zeta}$ are the camera extrinsic and intrinsic calibration parameters from the local mapper. Note that the 3D point and calibration are fixed for this optimization. The optimization is run up to four times, and after each step each residual is run through a $\chi^{2}$ check. If the residual fails the $\chi^{2}$ check, it is removed from the next round of optimization. At most 50% of residuals can be removed this way. This process robustifies against bad feature tracks or noisy 3D points.
@@
After the vision-only optimization, the velocity ${}^{A}\mathbf{v}_{I_{k}}$ and IMU biases $\mathbf{b}_{g,k}$, $\mathbf{b}_{a,k}$ are estimated in an inertial-only optimization (only preintegration measurements) with poses of the previous keyframe and current frame fixed. For the preintegration, the angular velocity is integrated using the quaternion integrator [38] and the acceleration is double integrated using fourth order Runge-Kutta. In the case where the IMU measurements (angular velocity or linear acceleration) exceed a saturation threshold, the corresponding measurement noise is inflated to robustify the system against IMU saturation. In particular, the inertial-only optimization tries to minimize the following cost function $\mathcal{C}_{io}$:
@@
which is the preintegration [39, 40, 41] cost function between the previous keyframe at time $t_{i}$ and the current frame at $t_{k}$. We have the state $\hat{\bar{\mathbf{x}}}_{k}=\begin{bmatrix}{}_{A}^{I_{k}}{\bar{q}}^{\top}&{}^{A}{\mathbf{p}}_{I_{k}}^{\top}&{}^{A}\hat{\mathbf{v}}_{I_{k}}^{\top}&\hat{\mathbf{b}}_{g,k}^{\top}&\hat{\mathbf{b}}_{a,k}^{\top}\end{bmatrix}^{\top}$ which is an inertial state (note that the pose is fixed), and similarly for the completely fixed keyframe state $\mathbf{x}_{i}$. The terms ${}^{i}\bm{\alpha}_{k}$ and ${}^{i}\bm{\beta}_{k}$ are the preintegrated measurements between $t_{i}$ and $t_{k}$ defined as [41]:
@@
$\mathbf{Q}_{ik}$ is the covariance matrix for the preintegration residual, which can be inflated as discussed if IMU saturation detected.
Note that in the above we have chosen to optimize motion by decoupling the visual and inertial measurements. This is because we found that tightly coupling the visual and inertial measurements in motion tracking (e.g., as in [5]) can sometimes pull the solution away from one that satisfies the reprojection error, as seen in Fig. 4, which creates wobblier pose output than desired for many applications. The problem is exacerbated when tracking a small number of features as in our system – since the vision terms can not outweigh the inertial – and the noisiness of the monocular depth does not help either. We emphasize that our decoupled approach is only possible with nonlinear optimization and not with filtering, as filters naturally tightly couple the measurements by using inertial measurements for propagation and visual ones for update, which adds priors from both measurements to the pose states.